OpenStudy (anonymous):

A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 2 experience points. On level 2, a player must earn 4 experience points. On level 3, a player must earn 7 experience points. On level 4, a player must earn 11 experience points. How many experience points must a player earn on level 7?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

@vzfreakz help

OpenStudy (anonymous):

There's a pattern into this, well an uneven pattern. You can tell from looking at 2, 4, 7, 11. It goes by 2, 3, & 4. Do you get where I'm going with this?

OpenStudy (anonymous):

@gislhaine1

OpenStudy (anonymous):

For level \(n\), the experience required \(a_n\) is \[\large a_n=n+a_{n-1}\]

OpenStudy (anonymous):

If you keep repeating that pattern, you'll get your answer.

OpenStudy (anonymous):

its really confusing

OpenStudy (anonymous):

Did you see the pattern that I done for you above?

OpenStudy (anonymous):

yes but i got 22

OpenStudy (anonymous):

What's 2+2? 4 4+3? 7 7+4? 11 Okay, do this - what's 11+5 =?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

thk you so much

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

Correct, 15 experience points that has been givin'. Then your next step would to add 15 to 6. What do you get for that?

OpenStudy (anonymous):

21

OpenStudy (anonymous):

It would be 21. Now it was asking for "How many experience points must a player earn on level 7?" Add 21 to 7 then that's your answer.

OpenStudy (anonymous):

aww thk you can u help me with another one

OpenStudy (anonymous):

11+5=16

OpenStudy (anonymous):

Sure. I'll try. But it would be better off if you ask this on a new open question. Not here.

OpenStudy (anonymous):

Oops. Must have been mis-calculated lol @SithsAndGiggles

OpenStudy (anonymous):

i got 28 its not on the answer choice

OpenStudy (anonymous):

Add 16 + 6

OpenStudy (anonymous):

It would be 29. I hoping that's the answer.

OpenStudy (anonymous):

Alternative method: \[\large \begin{align*}a_7&=7+a_6\\ &=7+6+a_5\\ &=7+6+5+a_4\\ &=7+6+5+11\\ &=29 \end{align*}\]

OpenStudy (anonymous):

Still the same answer, but yours was the easy way.